# On the Optimal Time Complexities in Decentralized Stochastic Asynchronous Optimization

### v1

@@ 1 Introduction
where $f\,:\,\mathbb{R}^{d}\times\mathbb{S}_{\xi}\rightarrow\mathbb{R},$ and $\mathcal{D}_{\xi}$ is a distribution on a non-empty set $\mathbb{S}_{\xi}.$ For a given $\varepsilon>0,$ we want to find a possibly random point $\bar{x},$ called an $\varepsilon$–stationary point, such that ${\mathbb{E}}[\left\|\nabla f(\bar{x})\right\|^{2}]\leq\varepsilon.$ We analyze the heterogeneous setup and the convex setup with smooth and non-smooth functions in Sections B and C.
@@ 2.2 Parallel optimization without communication costs
which depends on the time $\max_{i\in[n]}h_{i}$ of the slowest worker. There are many other more advanced methods including Picky SGD (Cohen et al.,, 2021), Asynchronous SGD (e.g., (Mishchenko et al.,, 2022; Koloskova et al.,, 2022)), and Rennala SGD (Tyurin and Richtárik,, 2023) that are designed to be robust to workers’ chaotic computation times. Under the assumption that the computation times of the workers are heterogeneous and bounded by $\{h_{i}\},$ Tyurin and Richtárik, (2023) showed that Rennala SGD is the first method that achieves the optimal time complexity
@@ 2.3 Parallel optimization with communication costs $\tau_{i\to j}$
We now consider the setup where workers’ communication times can not be ignored. This problem leads to a research field called decentralized optimization. This setup is the primary case for us. All $n$ workers calculate stochastic gradients in parallel and communicate with each other. Numerous works consider this setup, and we refer to Yang et al., (2019); Koloskova, (2024) for detailed surveys. Typically, methods in this setting use the gossip matrix framework (Duchi et al.,, 2011; Shi et al.,, 2015; Koloskova et al.,, 2021) and get an iteration converge rate that depends on the spectral gap of a mixing matrix. However, such rates do not give the physical time complexity of algorithms.
@@ 3 Contributions
We consider the class of functions that satisfy the setup and the assumptions from Section 1.1 and show that (informally) it is impossible to develop a method that will converge faster than (7) seconds. Next, we develop a new asynchronous stochastic method, Fragile SGD, that is nearly optimal (i.e., almost matches this lower bound; see Table 1 and Corollary 1). This is the first such method. It provably improves on Asynchronous SGD (Even et al.,, 2024) and all other synchronous and asynchronous methods (Bornstein et al.,, 2023). We also consider the heterogeneous setup (see Table 2 and Section B), where we discover the optimal time complexity by proving another lower bound and developing a new optimal method, Amelie SGD, with weak assumptions. The developed methods can guarantee the iteration complexity $\mathcal{O}\left(\nicefrac{{L\Delta}}{{\varepsilon}}\right)$ with arbitrarily heterogeneous random computation and communication times (Theorems 4 and 8). Our findings are extended to the convex setup in Section C, where we developed new accelerated methods, Accelerated Fragile SGD and Accelerated Amelie SGD.
@@ 4 Lower Bound
The intuition and meaning of the formula (8) is discussed in Section 5.2. Note that if we take $n=1$ and $\tau_{1\to 1}=0$ our lower bound reduces to the lower bound (3) up to a log factor. Moreover, if we take $n>1$ and $\tau_{i\to j}=0$ for all $i,j\in[n],$ then (8) reduces to (5) up to a log factor. Thus, (8) is nearly consistent with the lower bounds from (Arjevani et al.,, 2022; Tyurin and Richtárik,, 2023). We get an extra $\log n$ factor due to the generality of our setup. The reason is technical, and we explain it in Section D.5. In a nutshell, the lower problem reduces to the analysis of the concentration of the time series $y^{T}:=\min_{j\in[n]}y^{T}_{j}$ and $y^{T}_{j}:=\min_{i\in[n]}\left\{y^{T-1}_{i}+h_{i}\eta^{T}_{i}+\tau_{i\to j}\right\},$ where $y^{0}_{i}=0$ for all $i\in[n],$ and $\{\eta^{k}_{i}\}$ are i.i.d. geometric random variables. This analysis is not trivial due to the $\min_{i\in[n]}$ operations. Virtually all previous works that analyzed lower bounds did not have such a problem because they analyzed time series with a sum structure (e.g., $\bar{y}^{T}:=\bar{y}^{T-1}+\varrho^{T},$ where $\{\varrho^{k}\}$ are some random variables, and $\bar{y}^{0}=0$).
@@ 6 Example: Line or Circle
Let us consider Line graphs where we can get more explicit and interpretable formulas for (11). We analyze ND-Mesh, ND-Torus, and Star graphs in Section A. Surprisingly, even in some simple cases like Line or Star graphs, as far as we know, we provide new time complexity results and insights. In Section I, we show that our theoretical results are supported by computational experiments.
@@ A.1 ND-Mesh or ND-Torus
We now consider a generalization of Line graphs: ND-Mesh graphs. In Figures 3(a) and 3(b), we present examples of 2D-Mesh and 3D-Mesh. For simplicity, assume that $n=(2k+1)^{N}$ for some $k\in\mathbb{N}.$ The computation speeds $h_{i}=h$ for all $i\in[n],$ and the communicate speeds of the edges $\rho_{i\to j}=\rho$ if workers $i$ and $j$ are connected in a mesh and $\rho_{i\to j}=\infty$ for all other $i,j\in[n].$ Using geometrical reasoning, it is clear an index $j^{*},$ that minimizes (11), corresponds to the worker in the middle of a graph (13 in Figures 3(a) and 14 in Figures 3(b)). Therefore,
@@ A.3 General case
Appendix B Heterogeneous Setup
@@
B.1 Lower bound
In Section E, we prove the following lower bound.
@@ B.1 Lower bound
B.2 Amelie SGD: optimal method in the heterogeneous setting
@@ B.3 Discussion
B.4 Comparison with previous methods
@@
Appendix C Convex Functions in the Homogeneous and Heterogeneous Setups
@@
C.1 Assumptions in convex world
@@
C.2 Homogeneous setup and nonsmooth case
@@
The proof of Theorem 10 is almost the same as in Theorem 4 (see Section F). The proof of Theorem 4 states that the steps of Fragile SGD are equivalent to the classical SGD method. Thus, we can use the classical result from the literature (Lan,, 2020). Using Theorem 22, we get
@@
C.3 Homogeneous setup and smooth case
@@
C.4 Heterogeneous setup and nonsmooth case
@@
The proof of Theorem 14 is almost the same as in Theorems 8 (see Section G). The proof of Theorem 8 states that the steps of Amelie SGD are equivalent to the classical SGD method. Thus, we can use the classical result from the literature (Lan,, 2020). Using Theorem 22, we get
@@
C.5 Heterogeneous setup and smooth case
Consider the optimization problem (16). In this section, we use the same idea as in Section C.3. We will modify Amelie SGD and, instead of {NoHyper}Line LABEL:line:step_heter from Algorithm LABEL:alg:receiver_heter, we use the lines (20). We call such a method the Accelerated Amelie SGD method.
@@ C.7 Previous works
Appendix D Lower Bound: Diving Deeper into the Construction
@@ D.2 Lower bound
D.3 Proof sketch of Theorem 19
Let us provide a proof sketch that will give intuition behind the theorem. The full proof starts in Section D.4.
@@ D.3 Proof sketch of Theorem 19
The first part of the proof is standard (Carmon et al.,, 2020; Arjevani et al.,, 2022; Tyurin and Richtárik,, 2023; Huang et al.,, 2022; Lu and De Sa,, 2021), and we delegate it to Section D.4. For any algorithm, we construct oracles and a “worst-case” function such that
@@
D.4 Full proof of Theorem 19
@@
We fix $\lambda>0,$ $T\in\mathbb{N}$ and take the function $f(x):=\frac{L\lambda^{2}}{l_{1}}F_{T}\left(\frac{x}{\lambda}\right),$ where the function $F_{T}$ is defined in Section D.4. Note that the function $f$ is $L$-smooth:
@@
In the following lemma we use the notation of this theorem. We prove it in Section D.5.
@@ D.5 Proof of Lemma 2
Recall that $\bar{t}_{j}>0$ is a parameter. In the following technical lemma, we choose $\bar{t}_{j}$ and show that the second term in (44) is “small.” We prove it in Section D.6.
@@ D.6 Proof of Lemma 3
Appendix E Lower Bound in the Heterogeneous Setup
@@
Appendix F Proof of the Time Complexity for Homogeneous Case
@@
Appendix G Proof of the Time Complexity for Heterogeneous Case
@@
Appendix H Classical SGD Theory
@@ Appendix I Experiments
The results are presented in Figures 4, 5, and 6. The plots are fully consisted with Table 1. One can see that when the communication is fast (Fig. 4), there is no big difference between the methods because both Fragile SGD and Minibatch SGD use all workers in the optimization steps. However, when we start decreasing the communication speed, we observe that Fragile SGD converges faster. We looked deeper into the optimization processes of Fragile SGD in Figure 6 and observed that only 13 of 100 workers contribute to the optimization process for the batch size $S=120.$ Other workers are too far away from the pivot worker, and their contributions can only slow down optimization.

### v2

@@ 1 Introduction
where $f\,:\,\mathbb{R}^{d}\times\mathbb{S}_{\xi}\rightarrow\mathbb{R},$ and $\mathcal{D}_{\xi}$ is a distribution on a non-empty set $\mathbb{S}_{\xi}.$ For a given $\varepsilon>0,$ we want to find a possibly random point $\bar{x},$ called an $\varepsilon$–stationary point, such that ${\mathbb{E}}[\left\|\nabla f(\bar{x})\right\|^{2}]\leq\varepsilon.$ We analyze the heterogeneous setup and the convex setup with smooth and non-smooth functions in Sections C and D.
@@ 2.2 Parallel optimization without communication costs
which depends on the time $\max_{i\in[n]}h_{i}$ of the slowest worker. There are many other more advanced methods including Picky SGD (Cohen et al.,, 2021), Asynchronous SGD (e.g., (Recht et al.,, 2011; Nguyen et al.,, 2018; Mishchenko et al.,, 2022; Koloskova et al.,, 2022)), and Rennala SGD (Tyurin and Richtárik,, 2023) that are designed to be robust to workers’ chaotic computation times. Under the assumption that the computation times of the workers are heterogeneous and bounded by $\{h_{i}\},$ Tyurin and Richtárik, (2023) showed that Rennala SGD is the first method that achieves the optimal time complexity
@@ 2.3 Parallel optimization with communication costs $\tau_{i\to j}$
We now consider the setup where workers’ communication times can not be ignored. This problem leads to a research field called decentralized optimization. This setup is the primary case for us. All $n$ workers calculate stochastic gradients in parallel and communicate with each other. Numerous works consider this setup, and we refer to Yang et al., (2019); Koloskova, (2024) for detailed surveys. Typically, methods in this setting use the gossip matrix framework (Duchi et al.,, 2011; Shi et al.,, 2015; Koloskova et al.,, 2021) and get an iteration converge rate that depends on the spectral gap of a mixing matrix. However, such rates do not give the physical time of algorithms (see also Section B).
@@ 3 Contributions
We consider the class of functions that satisfy the setup and the assumptions from Section 1.1 and show that (informally) it is impossible to develop a method that will converge faster than (7) seconds. Next, we develop a new asynchronous stochastic method, Fragile SGD, that is nearly optimal (i.e., almost matches this lower bound; see Table 1 and Corollary 1). This is the first such method. It provably improves on Asynchronous SGD (Even et al.,, 2024) and all other synchronous and asynchronous methods (Bornstein et al.,, 2023). We also consider the heterogeneous setup (see Table 2 and Section C), where we discover the optimal time complexity by proving another lower bound and developing a new optimal method, Amelie SGD, with weak assumptions. The developed methods can guarantee the iteration complexity $\mathcal{O}\left(\nicefrac{{L\Delta}}{{\varepsilon}}\right)$ with arbitrarily heterogeneous random computation and communication times (Theorems 4 and 8). Our findings are extended to the convex setup in Section D, where we developed new accelerated methods, Accelerated Fragile SGD and Accelerated Amelie SGD.
@@ 4 Lower Bound
The intuition and meaning of the formula (8) is discussed in Section 5.2. Note that if we take $n=1$ and $\tau_{1\to 1}=0$ our lower bound reduces to the lower bound (3) up to a log factor. Moreover, if we take $n>1$ and $\tau_{i\to j}=0$ for all $i,j\in[n],$ then (8) reduces to (5) up to a log factor. Thus, (8) is nearly consistent with the lower bounds from (Arjevani et al.,, 2022; Tyurin and Richtárik,, 2023). We get an extra $\log n$ factor due to the generality of our setup. The reason is technical, and we explain it in Section E.5. In a nutshell, the lower problem reduces to the analysis of the concentration of the time series $y^{T}:=\min_{j\in[n]}y^{T}_{j}$ and $y^{T}_{j}:=\min_{i\in[n]}\left\{y^{T-1}_{i}+h_{i}\eta^{T}_{i}+\tau_{i\to j}\right\},$ where $y^{0}_{i}=0$ for all $i\in[n],$ and $\{\eta^{k}_{i}\}$ are i.i.d. geometric random variables. This analysis is not trivial due to the $\min_{i\in[n]}$ operations. Virtually all previous works that analyzed lower bounds did not have such a problem because they analyzed time series with a sum structure (e.g., $\bar{y}^{T}:=\bar{y}^{T-1}+\varrho^{T},$ where $\{\varrho^{k}\}$ are some random variables, and $\bar{y}^{0}=0$).
@@ 6 Example: Line or Circle
Let us consider Line graphs where we can get more explicit and interpretable formulas for (11). We analyze ND-Mesh, ND-Torus, and Star graphs in Section A. Surprisingly, even in some simple cases like Line or Star graphs, as far as we know, we provide new time complexity results and insights. In Section J, we show that our theoretical results are supported by computational experiments.
@@
7 Heterogeneous Setup
In Section C (in more details), we consider and analyze the problem
$\min\limits_{x\in\mathbb{R}^{d}}\Big{\{}f(x):=\frac{1}{n}\sum\limits_{i=1}^{n}{\rm\mathbb{E}}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}(x;\xi_{i})\right]\Big{\}},$
where $f_{i}\,:\,\mathbb{R}^{d}\times\mathbb{S}_{\xi_{i}}\rightarrow\mathbb{R}^{d}$ and $\xi_{i}$ are random variables with some distributions $\mathcal{D}_{i}$ on $\mathbb{S}_{\xi_{i}}.$ For all $i\in[n],$ worker $i$ can only access $f_{i}$. We show that the optimal time complexity is
$\textstyle\Theta\left(\frac{L\Delta}{\varepsilon}\max\left\{\max\limits_{i,j\in[n]}\mu_{i\to j},\max\limits_{i\in[n]}h_{i},\frac{\sigma^{2}}{n\varepsilon}\left(\frac{1}{n}\sum\limits_{i=1}^{n}h_{i}\right)\right\}\right)$ (16)
in the heterogeneous setting achieved by a new method, Amelie SGD (Algorithm LABEL:alg:alg_server_malenia). Amelie SGD is closely related to Fragile SGD but with essential algorithmic changes to make it work with heterogeneous functions. The obtained complexity (16) is worse than (12), which is expected because the heterogeneous setting is more challenging than the homogeneous setting.
8 Highlights of Experiments
In Section J, we present experiments with quadratic optimization problems, logistic regression, and a neural network to substantiate our theoretical findings. Here, we focus on highlighting the results from the logistic regression experiments:
Figure 3: The communication time $\rho=10$ seconds (Slow communication) in 2D-Mesh
On MNIST dataset (LeCun et al.,, 2010) with $100$ workers, Fragile SGD is much faster and has better test accuracy than Minibatch SGD.
Acknowledgments and Disclosure of Funding
The research reported in this publication was supported by funding from King Abdullah University of Science and Technology (KAUST): i) KAUST Baseline Research Scheme, ii) Center of Excellence for Generative AI, under award number 5940, iii) SDAIA-KAUST Center of Excellence in Artificial Intelligence and Data Science. The work of A.T. was partially supported by the Analytical center under the RF Government (subsidy agreement  [^00] D730321P5Q0002, Grant No. 70-2021-00145 02.11.2021).
@@ A.1 ND-Mesh or ND-Torus
We now consider a generalization of Line graphs: ND-Mesh graphs. In Figures 4(a) and 4(b), we present examples of 2D-Mesh and 3D-Mesh. For simplicity, assume that $n=(2k+1)^{N}$ for some $k\in\mathbb{N}.$ The computation speeds $h_{i}=h$ for all $i\in[n],$ and the communicate speeds of the edges $\rho_{i\to j}=\rho$ if workers $i$ and $j$ are connected in a mesh and $\rho_{i\to j}=\infty$ for all other $i,j\in[n].$ Using geometrical reasoning, it is clear an index $j^{*},$ that minimizes (11), corresponds to the worker in the middle of a graph (13 in Figures 4(a) and 14 in Figures 4(b)). Therefore,
@@ A.3 General case
Appendix B On the Connection to the Gossip Framework
Most of the previous methods were designed for a different setting, gossip-type communication. In fact, our setting is more general than the gossip communication. Indeed, recall that in the gossip communication, worker $i$ is allowed to get vectors from other workers through the operation $\sum_{j=1}^{n}w_{ij}x_{j},$ where $w_{ij}\in\{0,1\}$. This is equivalent to our setting for the case when the communication time $\rho_{ij}=\infty$ when $w_{ij}$ is zero, and $\rho_{ij}=1$ if $w_{ij}$ is not zero, and worker $i$ sums the received vectors. But our setting is richer since we allow different communication and computation times and allow workers to do whatever they like with vectors (not only to sum).
Moreover, the gossip framework codes the communication graph through the matrix $\{w_{ij}\}.$ We propose to code graphs through the times $\{\rho_{ij}\}$ ($\{\tau_{ij}\}$). Our approach is closer to real scenarios because, as we explained previously, it includes the gossip framework.
Appendix C Heterogeneous Setup
@@
C.1 Lower bound
In Section F, we prove the following lower bound.
@@ C.1 Lower bound
C.2 Amelie SGD: optimal method in the heterogeneous setting
@@ C.3 Discussion
C.4 Comparison with previous methods
@@
Appendix D Convex Functions in the Homogeneous and Heterogeneous Setups
@@
D.1 Assumptions in convex world
@@
D.2 Homogeneous setup and nonsmooth case
@@
The proof of Theorem 10 is almost the same as in Theorem 4 (see Section G). The proof of Theorem 4 states that the steps of Fragile SGD are equivalent to the classical SGD method. Thus, we can use the classical result from the literature (Lan,, 2020). Using Theorem 22, we get
@@
D.3 Homogeneous setup and smooth case
@@
D.4 Heterogeneous setup and nonsmooth case
@@
The proof of Theorem 14 is almost the same as in Theorems 8 (see Section H). The proof of Theorem 8 states that the steps of Amelie SGD are equivalent to the classical SGD method. Thus, we can use the classical result from the literature (Lan,, 2020). Using Theorem 22, we get
@@
D.5 Heterogeneous setup and smooth case
Consider the optimization problem (17). In this section, we use the same idea as in Section D.3. We will modify Amelie SGD and, instead of {NoHyper}Line LABEL:line:step_heter from Algorithm LABEL:alg:receiver_heter, we use the lines (21). We call such a method the Accelerated Amelie SGD method.
@@ D.7 Previous works
Appendix E Lower Bound: Diving Deeper into the Construction
@@ E.2 Lower bound
E.3 Proof sketch of Theorem 19
Let us provide a proof sketch that will give intuition behind the theorem. The full proof starts in Section E.4.
@@ E.3 Proof sketch of Theorem 19
The first part of the proof is standard (Carmon et al.,, 2020; Arjevani et al.,, 2022; Tyurin and Richtárik,, 2023; Huang et al.,, 2022; Lu and De Sa,, 2021), and we delegate it to Section E.4. For any algorithm, we construct oracles and a “worst-case” function such that
@@
E.4 Full proof of Theorem 19
@@
We fix $\lambda>0,$ $T\in\mathbb{N}$ and take the function $f(x):=\frac{L\lambda^{2}}{l_{1}}F_{T}\left(\frac{x}{\lambda}\right),$ where the function $F_{T}$ is defined in Section E.4. Note that the function $f$ is $L$-smooth:
@@
In the following lemma we use the notation of this theorem. We prove it in Section E.5.
@@ E.5 Proof of Lemma 2
Recall that $\bar{t}_{j}>0$ is a parameter. In the following technical lemma, we choose $\bar{t}_{j}$ and show that the second term in (45) is “small.” We prove it in Section E.6.
@@ E.6 Proof of Lemma 3
Appendix F Lower Bound in the Heterogeneous Setup
@@
Appendix G Proof of the Time Complexity for Homogeneous Case
@@
Appendix H Proof of the Time Complexity for Heterogeneous Case
@@
Appendix I Classical SGD Theory
@@
The results are presented in Figures 5, 6, and 7. The plots are fully consisted with Table 1. One can see that when the communication is fast (Fig. 5), there is no big difference between the methods because both Fragile SGD and Minibatch SGD use all workers in the optimization steps. However, when we start decreasing the communication speed, we observe that Fragile SGD converges faster. We looked deeper into the optimization processes of Fragile SGD in Figure 7 and observed that only 13 of 100 workers contribute to the optimization process for the batch size $S=120.$ Other workers are too far away from the pivot worker, and their contributions can only slow down optimization.